# A search using GEO600 for gravitational waves coincident with fast radio bursts from SGR 1935+2154

### v1

@@ 3.1 Simulated waveforms to quantify sensitivity
  • Sine-Gaussians: The majority of the simulated waveforms we use are Sine-Gaussians, which can model starquakes and certain neutron-star f-modes. They are described in Eq. 1 of Abbott et al. (2017). Most of these injections are performed with inclinations chosen randomly, but we also employ some optimally inclined (circular polarization only, emitted face-on to the observer) waveforms near the expected f-modes at $\sim$2000 Hz to better constrain our sensitivity to these models. In all the injected waveforms, we use a quality factor $Q=9$ following Abbott et al. (2021, 2023), with central frequencies $f_{0}$ spanning from 300 Hz to 3560 Hz, as shown in Table 2.
@@ 3.2 Long-duration search with PySTAMP
We use PySTAMP (Macquet et al., 2021) to target GW signals with durations longer than $1$ s around the three FRBs with coincident GEO600 data. The background distribution and the detection efficiency of the search are characterized using an off-source window that consists of $\sim 12$ hr of data centered on the event time, excluding the on-source window described above. The workflow of the pipeline is as follows. The data are first down-sampled from $16384$ Hz to $8192$ Hz, then high-pass filtered with a frequency cutoff of $40$ Hz to remove potential spectral leakage from lower frequencies. After these pre-processing steps, the resulting time series are split into $1$ s Hann-windowed segments with $50\%$ overlap. The fast Fourier transform is computed over each segment to build a time-frequency map (tf-map) with a resolution of $1$ s ×$\times 1$ Hz. For each frequency bin, the power spectral density (PSD) is estimated by taking the median of the squared modulus of the Fourier transform over $1320$ s of adjacent data (similar to Welch’s method but using the median instead of the mean), and a signal-to-noise ratio (S/N) tf-map is built by dividing the value of the Fourier transform in each pixel by the square root of the PSD. To identify candidate GW events, a pattern recognition algorithm is run over the tf-map. We use the burstegard algorithm (Prestegard, 2016) which identifies clusters of neighbouring pixels whose S/N is above a threshold of $2.5$. Clusters consisting of $5$ or more pixels are saved as candidate GW events. Each cluster is then assessed a ranking statistic $\Lambda$ that is the sum of the S/N of each of its pixels divided by the square root of the total number of pixels.
@@
Clusters found in the off-source window form the background of the search, and are used to estimate the false-alarm rate (FAR) of clusters found in the on-source window as a function of the ranking statistic $\Lambda$. PySTAMP is primarily intended to work on cross-correlated data from a pair of independent detectors, which allows for the simulation of an extended amount of background by shifting the time series of one detector with respect to the other. Such a method cannot be applied here in the single-detector GEO600 search. Hence, the background lifetime is limited to the duration of the off-source window, so the FAR of each cluster can only be estimated down to a minimum of $\sim 1$ per $12$ hours ($2.3\times 10^{-5}$ Hz). See Sec. 5 for further discussion of the limited FAR.
@@
Noise from GW detectors typically features narrowband spectral artifacts which appear as horizontal lines in a time-frequency representation. Because the PSD is estimated for each frequency bin by taking the median over neighbouring time segments, most of these lines are correctly factored into the PSD and do not generate high S/N pixels. However, we observe an excess of clusters in the off-source window around some specific frequencies, likely due to fluctuations of spectral lines around their central values. We therefore remove clusters for a narrow range of frequencies corresponding to known GEO600 spectral lines. In order to reject short, broadband noise transients (known as glitches), we also remove clusters for which more than $30\%$ of the total energy is contained within a single $1$ s time segment.
@@ 3.3 Short-duration search with X-Pipeline
We perform a search for short-duration unmodeled GW transients using X-Pipeline (Sutton et al., 2010; Was et al., 2012). While typically run as a coherent search across multiple detectors (such as in previous searches for GWs from FRBs (Abbott et al., 2023), gamma ray bursts (Abbott et al., 2021, 2022), and magnetars (Abbott et al., 2024)), we use X-Pipeline in a single-detector mode since GEO600 was the only GW detector collecting data at the time of the FRBs. X-Pipeline splits the PSD-whitened data into 64 s segments, then applies a Fourier-transform to produce time-frequency maps. The time-frequency pixels with amplitudes in the highest 1% that neighbor each other are clustered into candidate detection events. Each event is then assigned a ranking statistic based on the summed energy contained in the pixels. To determine the significance of these candidate events, we compare them against a distribution of background energies empirically measured in an identical manner from an “off-source” period. This is chosen to fall around (but not including) the time of the on-source data, and to be long enough to allow for meaningful significances to be calculated but not so long that the detector’s behavior is nonstationary. We employ a 24-hour off-source window, symmetric about each event’s time.
When X-Pipeline is run on data from multiple detectors as is typical, vetoes of problematic event candidates can be applied by utilizing the presumed coherence of any real GW event across detectors. This is unfortunately not an option in a single-detector search such as this, meaning that the search becomes more vulnerable to background noise. To improve the sensitivity of our search, we apply frequency-domain vetoes based on the distribution of time-frequency candidate events in the off-source window, for each FRB search. We veto narrow frequency bands ($\sim 10$ Hz bandwidth) where there is considerable excess noise; most of these vetoes corresponded to known spectral lines from the GEO600 detector.
@@ 4 Search results and limits on coincident emission
Figure 3: 90% upper limits on the emitted GW energy from SGR 1935+2154 coincident with FRBs, alongside GW energy limits from FRBs during the O3a observing run as reported in Abbott et al. (2023). We plot the short-duration ringdown (orange) and Sine-Gaussian (SG; blue) waveforms from Table 5, and the long-duration SG waveforms (aquamarine) from Table 3. The previous range of 90% limits from Abbott et al. (2023), based on the lower bounds of the 90% credible distance as reported in their Table A1, are plotted in vertical black lines. These are for a short-duration search; Abbott et al. (2023) did not perform a long-duration analysis.
Table 3: 50% and 90% upper limits on GW emission energy in ergs from the long-duration PySTAMP search. Each frequency corresponds to the central frequency $f_{0}$ of a Sine-Gaussian waveform with duration parameter equal to 10 s.
@@
Table 4: 50% upper limits on GW emission energy in ergs from the short-duration X-Pipeline search. Each injection waveform is defined as in Table 2. Dashes indicate that no limit was obtainable for this set of injections due to insufficient background or poor data quality.
@@
We do not find any candidate GW events in either the long-duration or short-duration compact-window searches, for any of the FRBs. For the long-duration PySTAMP search, no triggers survive the cuts described in Sec. 3.2 for the 2020 April and 2022 October FRBs. Five triggers survive for the 2020 October FRB, but the loudest trigger has a FAR of $\sim 1$ per 1000 s with a p-value of 0.76 and is thus not significant. In the short-duration X-Pipeline case, only the 2020 April, 2020 October, and 2022 October FRBs compact-window searches had enough background to be considered useful for GW searches, as described above. The only surviving trigger from these three searches is from the 2022 October FRB and has a p-value of 0.53, and thus is also not significant.
@@
For some injections in the short-duration search (mostly those at lower frequencies such as SG-D with $f_{0}=300$ Hz), limits could not be established since noise in the detector prevented sufficient recovery of the injected signals. Limited data availability and poor data quality around the time of the 2020 October 08 event meant that no injection reached 90% recovery, leading to the lack of 90% limits.
@@
The non-detection of GW emission from our analyzed FRBs implies that the GW-to-radio energy ratio must be less than $E_{\rm GW}/E_{\rm radio}\sim 8\times 10^{14}$ at the 90% level, for a time window from [-4, 4] s at a GW frequency of approximately 300 Hz. At the $\sim 2$ kHz frequencies close to the neutron star f-mode, $E_{\rm GW}/E_{\rm radio}\lesssim 1.7\times 10^{16}$ at the 90% level. The GW and radio energies for our analyzed FRBs are shown in Fig. 4, alongside the same quantities for FRBs from O3a analyzed in Abbott et al. (2023).
@@ 5 Discussion and conclusion
The previous best limits on coincident GW emission with FRBs were set by Abbott et al. (2023) using extragalactic FRBs observed during the O3a LVK observing run, using SG waveforms and X-Pipeline. Fig. 3 shows the 90% upper limits on the GW energy during our analyzed FRBs from SGR 1935+2154, compared previous limits presented as a range spanning the best and worst 90% limits from Abbott et al. (2023). In the short-duration search at approximately 300 Hz, the best previous 90% upper limit on GW energy was set at $3.4\times 10^{51}$ erg; at approximately 2 kHz, the best previous 90% upper limit was $7.9\times 10^{54}$ erg. We improve on the 300 Hz constraint by about two orders of magnitude, and the 2 kHz constraint by over four orders of magnitude. No previous long-duration searches around FRB events have been performed, so the long-duration search results presented here represent the first constraints on such emission. Studies have found that magnetar flares can emit up to $10^{48}-10^{49}$ erg in GW energy near the f-mode for $\sim 200$ ms (Ioka, 2001; Corsi & Owen, 2011)—a regime that is probed by our most stringent 50% short-duration constraints at approximately 2 kHz (see SG-H waveform in Table 4). We also slightly improve the upper limit on $E_{\rm GW}/E_{\rm radio}$, as shown in Fig. 4.
@@
We thank Teruaki Enoto, Chin-Ping Hu, and George Younes for assistance with their X-ray results. We also thank Vicky Kaspi, Kiyoshi Masui, and Kaitlyn Shin for helpful discussions.
@@
This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors also gratefully acknowledge the support of the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO 600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. The authors gratefully acknowledge the Italian Istituto Nazionale di Fisica Nucleare (INFN), the French Centre National de la Recherche Scientifique (CNRS) and the Netherlands Organization for Scientific Research (NWO), for the construction and operation of the Virgo detector and the creation and support of the EGO consortium. The authors also gratefully acknowledge research support from these agencies as well as by the Council of Scientific and Industrial Research of India, the Department of Science and Technology, India, the Science & Engineering Research Board (SERB), India, the Ministry of Human Resource Development, India, the Spanish Agencia Estatal de Investigación (AEI), the Spanish Ministerio de Ciencia, Innovación y Universidades, the European Union NextGenerationEU/PRTR (PRTR-C17.I1), the ICSC - CentroNazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by the European Union NextGenerationEU, the Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació i Transformació Digital with funds from the Tourist Stay Tax Law ITS 2017-006, the Conselleria d’Economia, Hisenda i Innovació, the FEDER Operational Program 2021-2027 of the Balearic Islands, the Conselleria d’Innovació, Universitats, Ciència i Societat Digital de la Generalitat Valenciana and the CERCA Programme Generalitat de Catalunya, Spain, the Polish National Agency for Academic Exchange, the National Science Centre of Poland and the European Union – European Regional Development Fund; the Foundation for Polish Science (FNP), the Polish Ministry of Science and Higher Education, the Swiss National Science Foundation (SNSF), the Russian Science Foundation, the European Commission, the European Social Funds (ESF), the European Regional Development Funds (ERDF), the Royal Society, the Scottish Funding Council, the Scottish Universities Physics Alliance, the Hungarian Scientific Research Fund (OTKA), the French Lyon Institute of Origins (LIO), the Belgian Fonds de la Recherche Scientifique (FRS-FNRS), Actions de Recherche Concertées (ARC) and Fonds Wetenschappelijk Onderzoek – Vlaanderen (FWO), Belgium, the Paris Île-de-France Region, the National Research, Development and Innovation Office Hungary (NKFIH), the National Research Foundation of Korea, the Natural Science and Engineering Research Council Canada, Canadian Foundation for Innovation (CFI), the Brazilian Ministry of Science, Technology, and Innovations, the International Center for Theoretical Physics South American Institute for Fundamental Research (ICTP-SAIFR), the Research Grants Council of Hong Kong, the National Natural Science Foundation of China (NSFC), the Leverhulme Trust, the Research Corporation, the National Science and Technology Council (NSTC), Taiwan, the United States Department of Energy, and the Kavli Foundation. The authors gratefully acknowledge the support of the NSF, STFC, INFN and CNRS for provision of computational resources.

### v2

@@ 3.1 Simulated waveforms to quantify sensitivity
  • Sine-Gaussians: The majority of the simulated waveforms we use are Sine-Gaussians, which can model starquakes and certain neutron-star f-modes. They are described in Eq. 1 of Abbott et al. (2017). Most of these injections are performed with inclinations chosen randomly, but we also employ some optimally inclined (circular polarization only, emitted face-on to the observer) waveforms near the expected f-modes at $\sim$2000 Hz to better constrain our sensitivity to these models. In all the injected waveforms, we use a quality factor $Q=9$ (the approximate number of cycles in the waveform) following Abbott et al. (2021, 2023), with central frequencies $f_{0}$ spanning from 300 Hz to 3560 Hz, as shown in Table 2.
@@ 3.2 Long-duration search with PySTAMP
We use PySTAMP (Macquet et al., 2021) to target GW signals with durations longer than $1$ s around the three FRBs with coincident GEO600 data. The background distribution and the detection efficiency of the search are characterized using an off-source window that consists of $\sim 12$ hr of data centered on the event time, excluding the on-source window described above. The workflow of the pipeline is as follows. The data are first down-sampled from $16384$ Hz to $8192$ Hz, and then high-pass filtered with a frequency cutoff of $40$ Hz to remove potential spectral leakage from lower frequencies. After these preprocessing steps, the resulting time series are split into $1$ s Hann-windowed segments with $50\%$ overlap. The fast Fourier transform is computed over each segment to build a time-frequency map (tf-map) with a resolution of $1$ s ×$\times 1$ Hz. For each frequency bin, the power spectral density (PSD) is estimated by taking the median of the squared modulus of the Fourier transform over $1320$ s of adjacent data (similar to Welch’s method but using the median instead of the mean), and a signal-to-noise ratio (S/N) tf-map is built by dividing the value of the Fourier transform in each pixel by the square root of the PSD. To identify candidate GW events, a pattern recognition algorithm is run over the tf-map. We use the burstegard algorithm (Prestegard, 2016) which identifies clusters of neighboring pixels whose S/N is above a threshold of $2.5$. Clusters consisting of $5$ or more pixels are saved as candidate GW events. Each cluster is then assessed a ranking statistic $\Lambda$ that is the sum of the S/N of each of its pixels divided by the square root of the total number of pixels.
@@
Clusters found in the off-source window form the background of the search and are used to estimate the false-alarm rate (FAR) of clusters found in the on-source window as a function of the ranking statistic $\Lambda$. PySTAMP is primarily intended to work on cross-correlated data from a pair of independent detectors, which allows for the simulation of an extended amount of background by shifting the time series of one detector with respect to the other. Such a method cannot be applied here in the single-detector GEO600 search. Hence, the background lifetime is limited to the duration of the off-source window, so the FAR of each cluster can only be estimated down to a minimum of $\sim 1$ per $12$ hours ($2.3\times 10^{-5}$ Hz). See Sec. 5 for further discussion of the limited FAR.
@@
Noise from GW detectors typically features narrowband spectral artifacts that appear as horizontal lines in a time-frequency representation (or vertical lines in the amplitude spectral density; see Fig. 1). Because the PSD is estimated for each frequency bin by taking the median over neighboring time segments, most of these lines are correctly factored into the PSD and do not generate high S/N pixels. However, we observe an excess of clusters in the off-source window around some specific frequencies, likely due to fluctuations of spectral lines around their central values. We therefore remove clusters for a narrow range of frequencies corresponding to known GEO600 spectral lines. In order to reject short, broadband noise transients (known as glitches), we also remove clusters for which more than $30\%$ of the total energy is contained within a single $1$ s time segment.
@@ 3.3 Short-duration search with X-Pipeline
We perform a search for short-duration unmodeled GW transients using X-Pipeline (Sutton et al., 2010; Was et al., 2012). While typically run as a coherent search across multiple detectors (such as in previous searches for GWs from FRBs (Abbott et al., 2023), gamma ray bursts (Abbott et al., 2021, 2022), and magnetars (Abbott et al., 2024)), we use X-Pipeline in a single-detector mode because GEO600 was the only GW detector collecting data at the time of the FRBs. X-Pipeline splits the PSD-whitened data into 64 s segments, then applies a Fourier transform to produce time-frequency maps. The time-frequency pixels with amplitudes in the highest 1% that neighbor each other are clustered into candidate detection events. Each event is then assigned a ranking statistic based on the summed energy contained in the pixels. To determine the significance of these candidate events, we compare them against a distribution of background energies empirically measured in an identical manner from an “off-source” period. This is chosen to fall around (but not including) the time of the on-source data and to be long enough to allow for meaningful significances to be calculated but not so long that the detector’s behavior is nonstationary. We employ a 24-hour off-source window, symmetric about each event’s time.
When X-Pipeline is run on data from multiple detectors as is typical, vetoes of problematic event candidates can be applied by utilizing the presumed coherence of any real GW event across detectors. This is unfortunately not an option in a single-detector search such as this, meaning that the search becomes more vulnerable to background noise. To improve the sensitivity of our search, we apply frequency-domain vetoes based on the distribution of time-frequency candidate events in the off-source window for each FRB search. We veto narrow frequency bands ($\sim 10$ Hz bandwidth) where there is considerable excess noise; most of these vetoes corresponded to known spectral lines from the GEO600 detector.
@@ 4 Search results and limits on coincident emission
Figure 3: The 90% upper limits on the emitted GW energy from SGR 1935+2154 coincident with FRBs, alongside GW energy limits from FRBs during the O3a observing run as reported in Abbott et al. (2023). We plot the short-duration ringdown (orange) and Sine-Gaussian (SG; blue) waveforms from Table 5, and the long-duration SG waveforms (aquamarine) from Table 3. The previous range of 90% limits from Abbott et al. (2023), based on the lower bounds of the 90% credible distance as reported in their Table A1, are plotted in vertical black lines. These are for a short-duration search; Abbott et al. (2023) did not perform a long-duration analysis.
Table 3: The 50% and 90% upper limits on GW emission energy in ergs from the long-duration PySTAMP search. Each frequency corresponds to the central frequency $f_{0}$ of a Sine-Gaussian waveform with duration parameter equal to 10 s.
@@
Table 4: The 50% upper limits on GW emission energy in ergs from the short-duration X-Pipeline search. Each injection waveform is defined as in Table 2. Dashes indicate that no limit was obtainable for this set of injections due to insufficient background or poor data quality.
@@
We do not find any candidate GW events in either the long-duration or short-duration compact-window searches for any of the FRBs. For the long-duration PySTAMP search, no triggers survive the cuts described in Sec. 3.2 for the 2020 April and 2022 October FRBs. Five triggers survive for the 2020 October FRB, but the loudest trigger has a FAR of $\sim 1$ per 1000 s with a p-value of 0.76 and is thus not significant. In the short-duration X-Pipeline case, only the 2020 April, 2020 October, and 2022 October FRBs compact-window searches had enough background to be considered useful for GW searches, as described above. The only surviving trigger from these three searches is from the 2022 October FRB and has a p-value of 0.53, and thus is also not significant.
@@
For some injections in the short-duration search (mostly those at lower frequencies such as SG-D with $f_{0}=300$ Hz), limits could not be established because noise in the detector prevented sufficient recovery of the injected signals. Limited data availability and poor data quality around the time of the 2020 October 08 event meant that no injection reached 90% recovery, leading to the lack of 90% limits.
@@
The nondetection of GW emission from our analyzed FRBs implies that the GW-to-radio energy ratio must be less than $E_{\rm GW}/E_{\rm radio}\sim 8\times 10^{14}$ at the 90% level, for a time window from [-4, 4] s at a GW frequency of approximately 300 Hz. At the $\sim 2$ kHz frequencies close to the neutron-star f-mode, $E_{\rm GW}/E_{\rm radio}\lesssim 1.7\times 10^{16}$ at the 90% level. The GW and radio energies for our analyzed FRBs are shown in Fig. 4, alongside the same quantities for FRBs from O3a analyzed in Abbott et al. (2023).
@@ 5 Discussion and conclusion
The previous best limits on coincident GW emission with FRBs were set by Abbott et al. (2023) using extragalactic FRBs observed during the O3a LVK observing run, using SG waveforms and X-Pipeline. Fig. 3 shows the 90% upper limits on the GW energy during our analyzed FRBs from SGR 1935+2154, compared previous limits presented as a range spanning the best and worst 90% limits from Abbott et al. (2023). In the short-duration search at approximately 300 Hz, the best previous 90% upper limit on GW energy was set at $3.4\times 10^{51}$ erg; at approximately 2 kHz, the best previous 90% upper limit was $7.9\times 10^{54}$ erg. We improve on the 300 Hz constraint by about two orders of magnitude, and the 2 kHz constraint by over four orders of magnitude. No previous long-duration searches around FRB events have been performed, so the long-duration search results presented here represent the first constraints on such emission. Studies have predicted that magnetar flares can emit up to $10^{48}-10^{49}$ erg in GW energy near the f-mode for $\sim 200$ ms (Ioka, 2001; Corsi & Owen, 2011)—a regime that is probed by our most stringent 50% short-duration constraints at approximately 2 kHz (see SG-H waveform in Table 4). We also slightly improve the upper limit on $E_{\rm GW}/E_{\rm radio}$, as shown in Fig. 4.
@@
We thank Teruaki Enoto, Chin-Ping Hu, and George Younes for assistance with their X-ray results. We also thank Vicky Kaspi, Kiyoshi Masui, and Kaitlyn Shin for helpful discussions, and the anonymous referee for their useful comments.
@@
This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation. The authors also gratefully acknowledge the support of the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO 600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. The authors gratefully acknowledge the Italian Istituto Nazionale di Fisica Nucleare (INFN), the French Centre National de la Recherche Scientifique (CNRS) and the Netherlands Organization for Scientific Research (NWO) for the construction and operation of the Virgo detector and the creation and support of the EGO consortium. The authors also gratefully acknowledge research support from these agencies as well as by the Council of Scientific and Industrial Research of India, the Department of Science and Technology, India, the Science & Engineering Research Board (SERB), India, the Ministry of Human Resource Development, India, the Spanish Agencia Estatal de Investigación (AEI), the Spanish Ministerio de Ciencia, Innovación y Universidades, the European Union NextGenerationEU/PRTR (PRTR-C17.I1), the ICSC - CentroNazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by the European Union NextGenerationEU, the Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació i Transformació Digital with funds from the Tourist Stay Tax Law ITS 2017-006, the Conselleria d’Economia, Hisenda i Innovació, the FEDER Operational Program 2021-2027 of the Balearic Islands, the Conselleria d’Innovació, Universitats, Ciència i Societat Digital de la Generalitat Valenciana and the CERCA Programme Generalitat de Catalunya, Spain, the Polish National Agency for Academic Exchange, the National Science Centre of Poland and the European Union – European Regional Development Fund; the Foundation for Polish Science (FNP), the Polish Ministry of Science and Higher Education, the Swiss National Science Foundation (SNSF), the Russian Science Foundation, the European Commission, the European Social Funds (ESF), the European Regional Development Funds (ERDF), the Royal Society, the Scottish Funding Council, the Scottish Universities Physics Alliance, the Hungarian Scientific Research Fund (OTKA), the French Lyon Institute of Origins (LIO), the Belgian Fonds de la Recherche Scientifique (FRS-FNRS), Actions de Recherche Concertées (ARC) and Fonds Wetenschappelijk Onderzoek – Vlaanderen (FWO), Belgium, the Paris Île-de-France Region, the National Research, Development and Innovation Office of Hungary (NKFIH), the National Research Foundation of Korea, the Natural Science and Engineering Research Council of Canada (NSERC), the Canadian Foundation for Innovation (CFI), the Brazilian Ministry of Science, Technology, and Innovations, the International Center for Theoretical Physics South American Institute for Fundamental Research (ICTP-SAIFR), the Research Grants Council of Hong Kong, the National Natural Science Foundation of China (NSFC), the Israel Science Foundation (ISF), the US-Israel Binational Science Fund (BSF), the Leverhulme Trust, the Research Corporation, the National Science and Technology Council (NSTC), Taiwan, the United States Department of Energy, and the Kavli Foundation. The authors gratefully acknowledge the support of the NSF, STFC, INFN and CNRS for provision of computational resources.